# A Survey of NL2SQL with Large Language Models: Where are we, and where are we going?

### v4

@@ II-D1 Rule-based Stage
In the early stages, statistical language models (e.g., semantic parsers) were used to interpret nl queries and convert them into sql queries using predefined rules [35, 36, 33, 37]. However, rule-based nl2sql methods face challenges in adaptability, scalability, and generalization. At this stage, natural language understanding was limited to the token level, with research primarily focused on single-table sql queries (see Figure 2(b)-①).
@@ II-D2 Neural Network-based Stage
To alleviate the limitations of rule-based methods, researchers explored neural networks for the nl2sql task. This led to the development of models based on sequence-to-sequence architectures and graph neural networks [38, 39, 40], which enhanced the handling of synonyms and intent understanding. Thus, research advanced from single-table scenarios to more complex multi-table scenarios (see Figure 2(b)-②). However, the generalization ability of these methods is still limited by model size and the availability of sufficient training data.
@@ II-D3 PLM-based Stage
The introduction of PLMs like BERT [41] and T5 [42] in 2018 led to significant advancements in nl2sql methods based on PLMs [43, 7, 44], achieving competitive performance on various benchmarks (see Figure 2(b)-③). At this stage, PLMs-based nl2sql models trained on large corpora have greatly enhanced natural language understanding, resolving approximately 80% of cases in the Spider dataset [45]. However, accuracy drops to about 50% on the extra hard cases of Spider [46]. In addition, these models still face challenges in handling complex schemas.
Remark: PLMs vs. LLMs Figure 3 shows the key differences between LLMs and PLMs. LLMs are a subset of PLMs, distinguished by their advanced language understanding and emergent capabilities [47, 48]. The emergent abilities allow LLMs to perform nl2sql tasks directly using prompts. In contrast, PLMs generally require additional pre-training or fine-tuning for acceptable nl2sql performance.
@@ II-D4 LLM-based Stage
LLMs demonstrate unique emergent capabilities that surpass traditional PLMs in NLP tasks, marking a new paradigm for nl2sql solutions. These LLM-based nl2sql methods have become the most representative solutions in the current nl2sql landscape [5, 49, 6, 50]. Current research focuses on optimizing prompt design [6] and fine-tuning LLMs [49]. For example, DAIL-SQL [6] utilizes the GPT-4 with effective prompt engineering techniques, achieving strong results on the Spider dataset [45]. Meanwhile, CodeS [49] builds an LLM specifically for nl2sql tasks by pretraining StarCoder [51] on a large nl2sql-related corpus, showing solid performance on benchmarks like BIRD [52]. At this stage, LLMs’ emergent capabilities have significantly improved natural language understanding, shifting the task’s focus toward database-specific challenges. New benchmarks like BIRD [52] and BULL [50] emphasize handling massive tables and domain-specific solutions (see Figure 2(b)-④).
NL2SQL Solutions in LLMs Era. Broadly speaking, there are two major approaches to leverage the capabilities of LLMs for nl2sql: 1) in-context learning, and 2) pre-train/fine-tune LLMs specialized for nl2sql.
In-Context Learning for NL2SQL. For in-context learning nl2sql methods, the goal is to optimize the prompt function $P$ to guide the LLMs, which can be formulated as follows:
@@
where k denotes additional information or domain-specific knowledge related to nl or db. $P$ is a prompt function that transforms the input (nl, db, k) into a suitable textual prompt for the LLMs. An appropriately designed $P$ can effectively guide the LLMs to perform the nl2sql task more accurately.
Employing in-context learning strategies for nl2sql treats LLMs as off-the-shelf tools, without modifying their parameters. However, if users have sufficient training data or hardware resources, calibrating the LLMs’ parameters can enhance performance and accuracy, allowing the model to be more closely tailored to the specific nl2sql task.
Pre-train and Fine-tune LLMs for NL2SQL. Fully optimizing the parameters of LLMs for nl2sql involves two key stages: pre-train and fine-tune, which can be formulated as follows:
@@
In the subsequent fine-tuning stage, the pre-trained model is further adjusted on a more specialized dataset $\mathcal{D}_{f}$, which is closely aligned with the nl2sql task. This targeted training refines the model’s capabilities, enabling it to more effectively interpret and generate sql based on natural language queries.
III Language Model-powered nl2sql Overview
Figure 4: An Overview of NL2SQL Modules in the LLM Era.
In this section, we first summarize the key modules of PLM- and LLM-powered nl2sql solutions, as shown in Figure 4. Next, we compare the key differences between the modules of existing nl2sql solutions in Table I.
Pre-Processing Methods. Pre-processing serves as an enhancement to the model’s inputs in the nl2sql parsing process. Although not strictly necessary, pre-processing significantly contributes to the refinement of nl2sql parsing [53].
  • Schema Linking: This key module identifies the most relevant tables and columns from nl2sql (Section IV-A).
@@ III Language Model-powered nl2sql Overview
NL2SQL Translation Methods. This is the core of nl2sql solution, responsible for converting input nl queries into sql.
@@
  • Task-specific Prompt Strategy: This module provides tailored guidance for the nl2sql model, optimizing the nl2sql translation workflow (Section V-C).
@@
  • N-best Rankers Strategy: It aims to rerank the top-$k$ results generated by the nl2sql model to enhance query accuracy (Section VI-D).
TABLE I: Comparisons of Existing NL2SQL Solutions.
Methods Years Finetuning Pre-Processing NL2SQL Translation Methods Post-Processing
@@
BRIDGE [79] 2020 - - Encoder-Only Sequential Encoding - - Others - - - -
IV Pre-Processing Strategies for NL2SQL
The pre-processing step is crucial in the nl2sql translation process, as it identifies relevant tables and columns (i.e., Schema Linking) and retrieves necessary database contents or cell values (i.e., DB Content Retrieval) to support sql query generation What’s more, it enriches context by incorporating domain-specific knowledge (i.e., Additional Information Acquisition), which can improve the understanding of the query context and correct errors to prevent their propagation.
@@ IV-A Schema Linking
Schema linking aims to identify the tables and columns relevant to the given nl query, ensuring accurate mapping and processing of key information within the limited input. This step is essential for improving the performance of the nl2sql task. In the LLM era, schema linking has become even more critical due to the input length limitations of LLMs.
@@ IV-A1 String Matching-based Schema Linking
Early research [80, 81, 39] primarily focused on string matching techniques for schema linking. These methods use similarity measures between the nl queries and schemas to identify relevant mappings. Both exact and approximate matching techniques are commonly employed in this approach.
Exact matching, as used by IRNet [80], identifies matches when candidates are identical or when one is a substring of the other. While this method can detect straightforward links, it may lead to false positives when candidates share common words. Approximate string matching techniques, like the Damerau–Levenshtein distance [82] used in ValueNet [83], help identify matches despite spelling variations or errors.
However, these methods struggle with handling synonyms and are not robust enough to manage vocabulary variations, limiting their effectiveness in complex nl2sql tasks.
@@ IV-A2 Neural Network-based Schema Linking
To alleviate the above limitations, researchers have employed deep neural networks to align database schemas with natural language queries [84, 53, 76, 7, 85]. These methods can better parse complex semantic relationships between nl queries and database structures. DAE [84] frames schema linking as a sequential tagging problem, using a two-stage anonymization model to capture semantic relationships between schema and nl. However, without an annotated corpus, DAE does not evaluate schema linking’s impact on nl2sql performance. To address this, SLSQL [53] annotates schema linking information in Spider dataset [45], enabling a systematic, data-driven study. RESDSQL [7] introduces a ranking-enhanced encoding framework for schema linking, using a cross-encoder to prioritize tables and columns based on classification probabilities. FinSQL [50] uses a parallel cross-encoder to retrieve relevant schema elements, significantly reducing linking time.
@@ IV-A3 In-Context Learning for Schema Linking
C3-SQL [62] employs zero-shot prompts with GPT-3.5 using self-consistency for table and column linking. For table linking, tables are ranked by relevance and listed; for column linking, columns are ranked within relevant tables and outputted as a dictionary, prioritizing matches with question terms or foreign keys. MAC-SQL [66] proposes a multi-agent collaborative framework for nl2sql, where the Selector agent handles schema linking, activated only when the database schema prompt exceeds a specified length. CHESS [54] utilizes GPT-4 to extract keywords from both nl and evidence (additional information from BIRD [52]), implementing a three-stage schema pruning protocol with different prompts.
@@ IV-B1 String Matching-based Methods
IRNet [80] uses n-grams, treating text between quotes as cell values. RESDSQL [7] uses the Longest Common Substring algorithm [88] to find the longest sequence shared between strings. BRIDGE [79] advances this with an anchor text matching technique that automatically extracts cell values from the nl. Using heuristics, it calculates the maximum sequence match to define matching boundaries, excluding irrelevant substrings and adjusts thresholds for accuracy.
@@ IV-B2 Neural Network-based Methods
TABERT [89] uses a method called database content snapshots to encode relevant database content for an nl query, employing attention mechanisms to manage information across cell value representations in different rows. Another approach leverages graph relationships to represent database content. For example, IRNet [80] uses the knowledge graph ConceptNet [90] to find and link relevant cell values, assigning types based on exact or partial matches. RAT-SQL [76] further enhances structural reasoning by modeling the relationship between cell values and the nl query, identifying column-value relationships where the query value is part of the column’s candidate cell value.
@@ IV-B3 Index Strategy for Database Content Retrieval
Efficiently retrieving relevant cell values is crucial for the performance of nl2sql systems, especially with large datasets. Indexing is a key method for improving retrieval efficiency by enabling faster access to relevant cell values [54, 49].
@@ IV-C Additional Information Acquisition
Additional information, such as domain knowledge, plays a crucial role in enhancing nl2sql models’ understanding of nl queries, schema linking, and overall nl2sql translation. This information can provide demonstration examples, domain knowledge, formulaic evidence, and format information for the nl2sql backbone model or specific modules, thereby enhancing the quality of generated results. We categorize existing strategies into the following two groups: 1) Sample-based Methods, and 2) Retrieval-based Methods.
@@ IV-C1 Sample-based Methods
With advancements in LLMs and in-context learning techniques, researchers often incorporate additional information into the textual inputs (i.e., prompts) alongside demonstration examples. DIN-SQL [5] integrates additional information through few-shot learning across multiple stages, such as schema linking, query classification, task decomposition, and self-correction. This enables DIN-SQL to handle challenges like complex schema links, multiple table joins, and nested queries. Other works [65, 49, 54] employ similar strategies. In real-world databases, cross-domain knowledge or evidence is often available. For example, BIRD [52] contains domain knowledge which is crucial for various nl2sql works [6, 55, 54, 49, 46].
NL2SQL Translation Methods Encoding Strategy V-A) Sequential Encoding Strategy C3-SQL [62], PET-SQL [57], CHESS [54], DAIL-SQL [6], DIN-SQL [5], CatSQL [67] , BRIDGE [79], MAC-SQL [66], TA-SQL [56], RESDSQL [7], DTS-SQL [55], CodeS [49], N-best List Rerankers [71] , T5+NatSQL+Token Preprocessing [63] Graph-based Encoding Strategy SQLformer [68], RASAT [72], SHiP [70], SmBoP [77], RaSaP [78] , S²SQL [75], Graphix-T5 [43], G³R [69], RATSQL [76] Separate Encoding Strategy SC-Prompt [1], TKK [74] Decoding Strategy V-B) Greedy Search-based Decoding Strategy DAIL-SQL [6], DIN-SQL [5], C3-SQL [62], PET-SQL [57], CHESS [54] , CodeS [49], SuperSQL [46], TA-SQL [56] , MAC-SQL [66], DTS-SQL [55], T5+NatSQL+Token Preprocessing [63] Beam Search-based Decoding Strategy CatSQL [67], ZeroNL2SQL [44], SQLformer [68], SC-Prompt [1], SmBoP [77] , RaSaP [78], G³R [69], RESDSQL [7], RATSQL [76] Constraint-aware Incremental Decoding Strategy TKK [74], PICARD [73], RASAT [72], SHiP [70], Graphix-T5 [43] , N-best List Rerankers [71] Other Decoding Strategies BRIDGE [79] Task-specific Prompt Strategy (§V-C) Chain-of-Thought CHESS [54], ACT-SQL [64], COE-SQL [58], TA-SQL [56], C3-SQL [62], G³R [69], MAC-SQL [66] Decomposition DTS-SQL [55], TKK [74], G³R [69], DEA-SQL [61], MAC-SQL [66], DIN-SQL [5] Intermediate Representation V-D) SQL-like Syntax Language Schema-free SQL [93], SyntaxSQLNet [94], SemQL [95], EditSQL [96], RAT-SQL [76], NatSQL [97], SHiP [70], QPL [96], RESDSQL [7], QDMR [98] SQL-like Sketch Structure SyntaxSQLNet [94], SC-prompt [1], CatSQL [67], ZeroNL2SQL [44], TA-SQL [56], RESDSQL [7]
Figure 5: A Taxonomy of NL2SQL Translation Methods based on their Design Choices.

### v5

@@ II-D1 Rule-based Stage
In the early stages, statistical language models (e.g., semantic parsers) were used to interpret nl queries and convert them into sql queries using predefined rules [35, 36, 32, 37]. However, rule-based Text-to-SQL methods face challenges in adaptability, scalability, and generalization. At this stage, natural language understanding was limited to the token level, with research primarily focused on single-table sql queries (see Figure 2(b)-①).
@@ II-D2 Neural Network-based Stage
To alleviate the limitations of rule-based methods, researchers explored neural networks for the Text-to-SQL task. This led to the development of models based on sequence-to-sequence architectures and graph neural networks [38, 39, 40], which enhanced the handling of synonyms and intent understanding. Thus, research advanced from single-table scenarios to more complex multi-table scenarios (see Figure 2(b)-②). However, the generalization ability of these methods is still limited by model size and the availability of sufficient training data.
@@ II-D3 PLM-based Stage
The introduction of PLMs like BERT [41] and T5 [42] in 2018 led to significant advancements in Text-to-SQL methods based on PLMs [43, 7, 44], achieving competitive performance on various benchmarks (see Figure 2(b)-③). At this stage, PLMs-based Text-to-SQL models trained on large corpora have greatly enhanced natural language understanding, resolving approximately 80% of cases in the Spider dataset [45]. However, accuracy drops to about 50% on the extra hard cases of Spider [46]. In addition, these models still face challenges in handling complex schemas.
Remark: PLMs vs. LLMs Figure 3 shows the key differences between LLMs and PLMs. LLMs are a subset of PLMs, distinguished by their advanced language understanding and emergent capabilities [47, 48]. The emergent abilities allow LLMs to perform Text-to-SQL tasks directly using prompts. In contrast, PLMs generally require additional pre-training or fine-tuning for acceptable Text-to-SQL performance.
@@ II-D4 LLM-based Stage
LLMs demonstrate unique emergent capabilities that surpass traditional PLMs in NLP tasks, marking a new paradigm for Text-to-SQL solutions. These LLM-based Text-to-SQL methods have become the most representative solutions in the current Text-to-SQL landscape [5, 49, 6, 50]. Current research focuses on optimizing prompt design [6] and fine-tuning LLMs [49]. For example, DAIL-SQL [6] utilizes the GPT-4 with effective prompt engineering techniques, achieving strong results on the Spider dataset [45]. Meanwhile, CodeS [49] builds an LLM specifically for Text-to-SQL tasks by pretraining StarCoder [51] on a large Text-to-SQL-related corpus, showing solid performance on benchmarks like BIRD [52]. At this stage, LLMs’ emergent capabilities have significantly improved natural language understanding, shifting the task’s focus toward database-specific challenges. New benchmarks like BIRD [52] and BULL [50] emphasize handling massive tables and domain-specific solutions (see Figure 2(b)-④).
Text-to-SQL Solutions in LLMs Era. Broadly speaking, there are two major approaches to leverage the capabilities of LLMs for Text-to-SQL: 1) in-context learning, and 2) pre-train/fine-tune LLMs specialized for Text-to-SQL.
In-Context Learning for Text-to-SQL. For in-context learning methods, the goal is to optimize the prompt function $P$ to guide the LLMs, which can be formulated as follows:
@@
where k denotes additional information or domain-specific knowledge related to nl or db. $P$ is a prompt function that transforms the input (nl, db, k) into a suitable textual prompt for the LLMs. An well-designed $P$ can effectively guide the LLMs to perform the Text-to-SQL task more accurately.
Employing in-context learning strategies for Text-to-SQL treats LLMs as off-the-shelf tools, without modifying their parameters. However, if users have sufficient training data or hardware resources, calibrating the LLMs’ parameters can enhance performance and accuracy, allowing the model to be more closely tailored to the specific Text-to-SQL task.
Pre-train and Fine-tune LLMs for Text-to-SQL. Fully optimizing the parameters of LLMs for Text-to-SQL involves two key stages, pre-train and fine-tune, formulated as follows:
@@
In the subsequent fine-tune stage, the pre-trained model is further adjusted on a more specialized dataset $\mathcal{D}_{f}$, which is closely aligned with the Text-to-SQL task. This targeted training refines the model’s capabilities, enabling it to more effectively interpret and generate sql based on nl queries.
III Language Model-powered Text-to-SQL Overview
Figure 4: An Overview of Text-to-SQL Modules in LLM Era.
While Text-to-SQL was initially designed as an end-to-end task, recent advances, particularly in the LLM era, have shifted towards a modular design. As shown in Figure 4, modern PLM- and LLM-based solutions typically decompose the task into three stages: Pre-Processing, Translation, and Post-Processing. Each stage contains specialized modules such as schema linking, intermediate representation, and execution-guided correction. This design reflects the increasing complexity of Text-to-SQL and aligns with the rising trend of multi-agent or multi-module collaboration. Table I further compares the key design choices across recent solutions.
Pre-Processing Methods. Pre-processing enhances inputs and plays a significant role in improving Text-to-SQL parsing [53].
  • Schema Linking: This module identifies the most relevant tables and columns from Text-to-SQL (Section IV-A).
@@ III Language Model-powered Text-to-SQL Overview
Translation Methods. This is the core of Text-to-SQL solution, responsible for converting input nl queries into sql.
@@
  • Task-specific Prompt Strategy: This module provides tailored guidance for the Text-to-SQL model, optimizing the Text-to-SQL translation workflow (Section V-C).
@@
  • N-best Rankers Strategy: It aims to rerank the top-$k$ results generated by the Text-to-SQL model to enhance query accuracy (Section VI-D).
Multi-agent Collaboration for Text-to-SQL in LLM Era. Building upon modular design principles, recent research has further introduced multi-agent architectures to tackle the Text-to-SQL task. In contrast to traditional monolithic systems, multi-agent frameworks assign specialized responsibilities to distinct agents, each dedicated to handling a specific subtask. This design facilitates enhanced division of labor and more effective coordination among components. For example, MAC-SQL [54] adopts a three-agent architecture, with separate agents for schema linking, query decomposition and generation, and execution-guided refinement. Similarly, CHASE-SQL [55] employs a divide-and-conquer approach, selecting relevant database content during preprocessing, generating SQL queries through multiple chain-of-thought pathways, and iteratively refining outputs through self-correction and ranking. Pushing the boundary further, Alpha-SQL [56] proposes a planning-centric autonomous agent framework that leverages LLMs in combination with Monte Carlo Tree Search (MCTS). This agent dynamically selects and activates the appropriate modules, such as schema linking and SQL generation, based on contextual reasoning and execution-based feedback. Alpha-SQL’s strategy-driven exploration and adaptive control offer robust generalization, avoiding the rigidity of pipeline-based approaches.
TABLE I: Comparisons of Existing Text-to-SQL Solutions.
Methods Years Finetuning Pre-Processing Text-to-SQL Translation Methods Post-Processing
@@
OpenSearch-SQL [57] 2025 - Decoder-Only Sequential Encoding Syntax Language COT Greedy Search -
CHASE-SQL [55] 2025 - Decoder-Only Sequential Encoding - Multi-COT Greedy Search -
ROUTE [58] 2025 Decoder-Only Sequential Encoding - - Greedy Search -
Alpha-SQL [56] 2025 - Decoder-Only Sequential Encoding - COT Greedy Search -
@@
BRIDGE [83] 2020 - - Encoder-Only Sequential Encoding - - Constraint-aware Incremental - - - -
IV Pre-Processing Strategies for Text-to-SQL
The pre-processing step is crucial in the Text-to-SQL translation process, as it identifies relevant tables and columns (i.e., Schema Linking) and retrieves necessary database contents or cell values (i.e., DB Content Retrieval) to support sql query generation What’s more, it enriches context by incorporating domain-specific knowledge (i.e., Additional Information Acquisition), which can improve the understanding of the query context and correct errors to prevent their propagation.
@@ IV-A Schema Linking
Schema linking aims to identify the tables and columns relevant to the given nl query, ensuring accurate mapping and processing of key information within the limited input. This step is essential for improving the performance of the Text-to-SQL task. In the LLM era, schema linking has become even more critical due to the input length limitations of LLMs.
@@ IV-A1 String Matching-based Schema Linking
Early research [84, 85, 39] primarily focused on string matching techniques for schema linking. These methods use similarity measures between the nl queries and db schemas to identify relevant mappings. IRNet [84] adopts exact matching, identifying links when candidates are identical or one is a substring of the other. While effective for simple cases, it may yield false positives due to shared words. To handle spelling variations, ValueNet [86] applies approximate matching via the Damerau–Levenshtein distance [87].
However, these methods struggle with handling synonyms and are not robust enough to manage vocabulary variations, limiting their effectiveness in complex Text-to-SQL tasks.
@@ IV-A2 Neural Network-based Schema Linking
To alleviate the above limitations, researchers have employed deep neural networks to align database schemas with natural language queries [53, 7]. These methods can better parse complex semantic relationships between nl queries and database schema.
DAE [88] frames schema linking as a sequential tagging problem, using a two-stage anonymization model to capture semantic relationships between schema and nl. SLSQL [53] annotates schema linking information in Spider dataset [45], enabling a systematic, data-driven study. RESDSQL [7] introduces a ranking-enhanced encoding framework for schema linking, using a cross-encoder to prioritize tables and columns based on classification probabilities. FinSQL [50] uses a parallel cross-encoder to retrieve relevant schema elements, significantly reducing linking time.
@@ IV-A3 In-Context Learning for Schema Linking
C3-SQL [67] employs zero-shot prompts with GPT-3.5 using self-consistency for table and column linking. For table linking, tables are ranked by relevance and listed; for column linking, columns are ranked within relevant tables and outputted as a dictionary, prioritizing matches with question terms or foreign keys. MAC-SQL [54] proposes a multi-agent collaborative framework for Text-to-SQL, where the Selector agent handles schema linking, activated only when the database schema prompt exceeds a specified length. CHESS [59] utilizes GPT-4 to extract keywords from both nl and evidence (additional information from BIRD [52]), implementing a three-stage schema pruning protocol with different prompts.
@@ IV-B1 String Matching-based Methods
IRNet [84] uses n-grams, treating text between quotes as cell values. BRIDGE [83] advances this with an anchor text matching technique that automatically extracts cell values from the nl. Using heuristics, it calculates the maximum sequence match to define matching boundaries, excluding irrelevant substrings and adjusts thresholds for accuracy.
@@ IV-B2 Neural Network-based Methods
TABERT [91] uses a method called database content snapshots to encode relevant database content for snl query, employing attention mechanisms to manage information across cell value representations in different rows. Another approach leverages graph relationships to represent database content. IRNet [84] uses the knowledge graph ConceptNet [92] to find and link relevant cell values, assigning types based on exact or partial matches. RAT-SQL [80] further enhances structural reasoning by modeling the relationship between cell values and the nl query, identifying column-value relationships where the query value is part of the column’s candidate cell value.
@@ IV-B3 Index Strategy for Database Content Retrieval
Efficiently retrieving relevant cell values is crucial for the performance of Text-to-SQL systems, especially with large datasets. Indexing is a key method for improving retrieval efficiency by enabling faster access to relevant cell values [59, 49].
@@ IV-C Additional Information Acquisition
Additional information, such as domain knowledge, plays a crucial role in enhancing Text-to-SQL models’ understanding of nl queries, schema linking, and overall Text-to-SQL translation. This information can provide demonstration examples, domain knowledge, formulaic evidence, and format information for the Text-to-SQL backbone model or specific modules, thereby enhancing the quality of generated results. We categorize existing strategies into the following two groups: 1) Sample-based Methods, and 2) Retrieval-based Methods.
@@ IV-C1 Sample-based Methods
With advancements in LLMs and in-context learning techniques, researchers often incorporate additional information into the textual inputs (i.e., prompts) alongside demonstration examples [70, 49, 59]. DIN-SQL [5] integrates additional information through few-shot learning across multiple stages. This helps DIN-SQL to handle challenges like complex schema links, multiple table joins, and nested queries. In practice, real-world databases often contain rich cross-domain knowledge that can serve as external evidence to support query generation. For example, BIRD [52] contains domain knowledge which is crucial for various Text-to-SQL works [6, 59]. Recent works [70, 49] also encode schema metadata(e.g., data types) into natural language to enhance context understanding.
Text-to-SQL Translation Methods Encoding Strategy V-A) Sequential Encoding C3-SQL [67], PET-SQL [62], CHESS [59], DAIL-SQL [6], DIN-SQL [5], CatSQL [71] , BRIDGE [83], MAC-SQL [54], TA-SQL [61], RESDSQL [7], CodeS [49], N-best List Rerankers [75] , T5+NatSQL+Token Preprocessing [68] Graph-based Encoding SQLformer [72], RASAT [76], SHiP [74], SmBoP [81], RaSaP [82] , S²SQL [79], Graphix-T5 [43], G³R [73], RATSQL [80] Separate Encoding SC-Prompt [1], TKK [78] Decoding Strategy V-B) Greedy Search-based Decoding DAIL-SQL [6], DIN-SQL [5], C3-SQL [67], PET-SQL [62], CHESS [59] , CodeS [49], SuperSQL [46], TA-SQL [61] , MAC-SQL [54], T5+NatSQL+Token Preprocessing [68] Beam Search-based Decoding CatSQL [71], ZeroNL2SQL [44], SQLformer [72], SC-Prompt [1], SmBoP [81] , RaSaP [82], G³R [73], RESDSQL [7], RATSQL [80] Constraint-aware Incremental Decoding TKK [78], PICARD [77], RASAT [76], SHiP [74], Graphix-T5 [43] , BRIDGE [83], N-best List Rerankers [75] Task-specific Prompt Strategy (§V-C) Chain-of-Thought CHESS [59], ACT-SQL [69], COE-SQL [63], TA-SQL [61], C3-SQL [67], G³R [73], MAC-SQL [54] Decomposition TKK [78], G³R [73], DEA-SQL [66], MAC-SQL [54], DIN-SQL [5] Intermediate Representation V-D) SQL-like Syntax Language Schema-free SQL [96], SyntaxSQLNet [97], SemQL [98], EditSQL [99], RAT-SQL [80], NatSQL [100], SHiP [74], QPL [99], RESDSQL [7], QDMR [101], OpenSearch-SQL [57] SQL-like Sketch Structure SyntaxSQLNet [97], SC-prompt [1], CatSQL [71], ZeroNL2SQL [44], TA-SQL [61], RESDSQL [7]
Figure 5: A Taxonomy of Text-to-SQL Translation Methods based on their Design Choices.